Determine the largest capacity of a cylindrical tank if its surface area (without lid) should be equal to S


1
Expert's answer
2021-10-18T16:25:53-0400

The surface area of the cylinder without the lid is "S=2\pi rh+\pi r^2" and the maximum volume will be "V=\pi r^2 h". From the first expression, we find "h=\cfrac{S}{2\pi r}-\cfrac{r}{2}" and we substitute this on the formula for S to find:


"V=\pi r^2 (\frac{S}{2\pi r}-\frac{r}{2} )=\cfrac{S}{2}\cdot r-\cfrac{\pi}{2}\cdot r^3"


We proceed to find the maximum for V in terms of r when V'(r) = 0:


"V'_{(r)}=\cfrac{d}{dr} \Bigg( \cfrac{S}{2}\cdot r-\cfrac{\pi}{2}\cdot r^3 \Bigg)
\\ V'_{(r)}= \cfrac{S}{2}-\cfrac{3\pi}{2}\cdot r^2=0
\\ \therefore V''_{(r)}= -3\pi\cdot r
\\ V'_{(r)}= \cfrac{S}{2}-\cfrac{3\pi}{2}\cdot r^2=0
\\ \cfrac{S}{2}=\cfrac{3\pi}{2}\cdot r^2 \implies r=\sqrt{\cfrac{S}{3\pi}}"


Once we find out the value for r, we proceed to substitute and find

"h= \cfrac{S}{2\pi r}-\cfrac{r}{2} = \cfrac{S}{2\pi \sqrt{\cfrac{S}{3\pi}}}-\cfrac{\sqrt{\cfrac{S}{3\pi}}}{2} 
\\ \therefore h=\cfrac{1}{2} \Bigg( \sqrt{\cfrac{3S}{\pi}}-\sqrt{\cfrac{S}{3\pi}} \, \Bigg)=\sqrt{\cfrac{S}{3\pi}} \Bigg( \cfrac{3-\sqrt{3}}{2} \, \Bigg)"


The fist result is valid because V''(r) is negative and this occurs when V(r) is at a maximum. In conclusion, the dimensions for the cylinder to have the maximum volume (with a surface area S) are "h=\sqrt{\cfrac{S}{3\pi}} \Bigg( \cfrac{3-\sqrt{3}}{2} \, \Bigg)" and "r=\sqrt{\cfrac{S}{3\pi}}".



Reference:

  • Thomas, G. B., & Finney, R. L. (1961). Calculus. Addison-Wesley Publishing Company.